# PIPA: A Unified Evaluation Protocol for Diagnosing Interactive Planning Agents

### v1

@@ Abstract
The growing capabilities of large language models (LLMs) in instruction-following and context-understanding lead to the era of agents with numerous applications. Among these, task planning agents have become especially prominent in realistic scenarios involving complex internal pipelines, such as context understanding, tool management, and response generation. However, existing benchmarks predominantly evaluate agent performance based on task completion as a proxy for overall effectiveness. We hypothesize that merely improving task completion is misaligned with maximizing user satisfaction, as users interact with the entire agentic process and not only the end result. To address this gap, we propose PIPA, a unified evaluation protocol that conceptualizes the behavioral process of interactive task planning agents within a partially observable Markov Decision Process (POMDP) paradigm. The proposed protocol offers a comprehensive assessment of agent performance through a set of atomic evaluation criteria, allowing researchers and practitioners to diagnose specific strengths and weaknesses within the agent’s decision-making pipeline. Our analyses show that agents excel in different behavioral stages, with user satisfaction shaped by both outcomes and intermediate behaviors. We also highlight future directions, including systems that leverage multiple agents and the limitations of user simulators in task planning.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed in real-world applications, primarily due to their ability to understand complex goals and devise structured sequences of action: a process known as “planning” (Wang et al., 2024). To assess and refine planning skills, researchers have introduced diverse benchmarks in web-based (Yao et al., 2022; Xie et al., 2024b), mobile (Deng et al., 2024), embodied (Choi et al., 2024), and automated testbeds (Park et al., 2023).
One of the most impactful planning applications is task planning, where agents generate and execute domain-specific plans (e.g., itineraries) to help users achieve goals while adhering to contextual constraints (Budzianowski et al., 2018; Rastogi et al., 2020; Xie et al., 2024a). While agents must accurately interpret user requests, leverage predefined tools, and engage in personalized dialogues, many benchmarks focus solely on final task completion, overlooking the agents’ intermediate behaviors and planning steps. Given that overall user satisfaction is shaped by multiple factors (e.g., efficiency and effectiveness) throughout prolonged interactions (Ashfaq et al., 2020), a narrow focus on outcome-based metrics can misrepresent an agent’s true effectiveness, especially in complex task-oriented applications (Cao et al., 2025; Xia et al., 2025; Acikgoz et al., 2025).
In addition, existing benchmarks across different domains often introduce bespoke evaluation frameworks and metrics, making it possible for the same agent to be evaluated using entirely different criteria. This fragmentation introduces inconsistencies in how agent behavior is interpreted and makes it difficult to draw generalizable conclusions about agent capabilities. Without a unified evaluation protocol that captures key aspects of agent behavior beyond task-specific outputs, it becomes challenging to meaningfully compare systems, track progress over time, and identify fundamental limitations in agent design.
To this end, we propose PIPA, a unified evaluation Protocol for Interactive Planning Agents. To generalize the evaluation criteria across arbitrary agentic scenarios, PIPA is formulated within the framework of a partially observed Markov Decision Process (POMDP), recognizing that real-world states (e.g., user intentions and satisfactions) are often hidden or only partially observable. By capturing these dynamics, PIPA defines domain- and task-agnostic evaluation criteria that can be easily instantiated through a set of atomic and easily measurable metrics. As illustrated in Figure 1, PIPA provides the following three key advantages:
@@
    Unified Evaluation: PIPA establishes a task-agnostic protocol for interactive task planning, enabling a consistent comparison across diverse benchmarks under a shared set of principles, which has been discussed for decades (Walker et al., 1997) but still has not been fully resolved.
@@
    Multi-Axis Diagnosis: By aligning with the POMDP framework, PIPA supports evaluation along multiple stages of the interaction pipeline, encompassing both intermediate decisions and final outcomes. This allows for a fine-grained diagnostic analysis of how different agents behave and impact performance.
@@
    Cross-Benchmark Comparisons: PIPA enables systematic comparisons across heterogeneous tasks and environments, thereby uncovering planning strategies and revealing performance trade-offs that may not be observable within the scope of individual benchmarks.
Figure 1: PIPA provides domain- and task-agnostic diagnostic evaluation protocols for interactive task agents following the POMDP paradigm.
Through extensive experiments, we demonstrate that different models exhibit distinct strengths and weaknesses within decision-making pipelines. Furthermore, human studies indicate that stage-specific evaluation using PIPA correlates more strongly with improvements in user satisfaction than with the final task completion metric alone. Lastly, analyses on combining agents during deployment and the reliability of user simulators highlight promising research directions for future scholars and practitioners.
@@ 2 Background: Bespoke Evaluation of Planning Agents
Recent benchmarks for evaluating agentic tasks highlight a variety of metrics tailored to capture different facets of agent performance, reflecting the complexity and multi-dimensional nature of interactive planning tasks, as illustrated in Table 5 in Appendix A. For instance, AgentBench (Liu et al., 2024) incorporates domain-specific metrics such as Success Rate, Win Rate, F1 Score, and Exact Match to assess agents comprehensively. Similarly, WebLINX (Lù et al., 2024) introduces granular evaluation dimensions, including Intent Match, a boolean indicator of correct intent prediction; Element Similarity, verifying correct function arguments; and Text Similarity, focusing on the lexical consistency of arguments. FlowBench (Xiao et al., 2024) further enhances the evaluation landscape by adding metrics focused explicitly on tool usage: Tool Invocation, measured by precision, recall, and F1 score for identifying the correct tool configurations; Success Rate, denoting the proportion of entirely successful sessions; and Task Progress, capturing the percentage of goals completed within a session. Meanwhile, $\tau$-Bench (Yao et al., 2025) utilizes more direct measures of success with Pass@k, a boolean indicator reflecting success across k attempts, and Pass^k, which requires successful outcomes for all k trials. Unlike PIPA, all such works tailor their evaluation metrics to their usage, thus lacking in a generalizable evaluation across benchmarks. Furthermore, the primary focus of most benchmarks is measuring task completion rates, which has been pointed out for several years (Ribeiro et al., 2020). Even for complex and subjective tasks such as information aggregation, the evaluation is done with final task success (Reddy et al., 2024; Yoran et al., 2024).
3 Proposed Evaluation Protocol: PIPA
@@
3.1 Components of PIPA
Metric Design Criteria
Figure 2: Decision theory-based taxonomy of evaluation metrics in PIPA. As mentioned in Walker et al. (1997), it should be cautious to generalize the original taxonomy to different agents and tasks. Following this, we provide distinct interpretations of each element, which will be described in further sections.
Interactive planning tasks are generally modeled using the POMDP paradigm, where agents devise plans based on partial observational data (Sun et al., 2024; He et al., 2024; Xie et al., 2024a; Yao et al., 2025). Building on the framework of Williams & Young (2007), we present an abstract pipeline consisting of $\mathcal{S}$ (a set of agent states), $\mathcal{A}$ (a set of actions), $\mathcal{O}$ (a set of observations) and $\mathcal{R}$ (a set of rewards) [^1] Our formulation adopts discrete spaces while acknowledging the existence of continuous concepts, such as transition and observation probabilities., along with an optional policy definition $\mathcal{P}$ (a set of global policies). The metrics of PIPA are defined for each phase within this pipeline, as various factors contribute to user satisfaction beyond a singular overall evaluation (Ashfaq et al., 2020; Siro et al., 2022). Consequently, PIPA aims to provide a comprehensive evaluation of agent performance, not just in terms of the overall results but also by diagnosing which specific phases may need improvement. Considering the fact that most behavioral stages of LLM agents are pipelined with natural languages, we employ a set of atomic LLM-based evaluations by following previous observations where atomic and decomposed evaluation criteria yield consistent and reliable outcomes (Liu et al., 2023; Saad-Falcon et al., 2024; Lee et al., 2025).
Furthermore, we align PIPA with the decision theory-based taxonomy (Walker et al., 1997), as illustrated in Figure 2. In line with this taxonomy, the evaluation metrics in PIPA (i.e., state consistency, tool efficiency, observation alignment, policy adherence, and task completion) contain a “disparate set of performance measures” that can collectively assess the fundamental objective: maximizing user satisfaction. By establishing a principled connection between decision theory and POMDP-based PIPA, we ground our evaluation framework in a principled model of sequential decision making under uncertainty. Specifically, we leverage the fact that POMDPs are a natural instantiation of decision theory, where the agent selects actions to maximize expected utility given a belief over latent states. This integration enables PIPA to measure user satisfaction holistically while remaining grounded in theoretical rigor.
Throughout this work, as visualized in Figure 1, we formulate the basic components of multi-turn, multi-step agents as follows: $T$ denotes the total number of turns (i.e., alternating utterances between the user and the agent) in an interactive session, $M_{t}$ represents the number of steps in the agent’s internal reasoning process within a single utterance at turn $t$, and $\{U_{\tau},A_{\tau}\}_{\tau=1}^{t}$ signifies the sequence of user inputs and agent responses observed up to turn $t$.
@@ 3.1.1 State Consistency ($\mathcal{S}$)
The intermediate step of the agent is pivotal in summarizing user requests and determining optimal subsequent actions (Shin et al., 2022; Wei et al., 2022). Given that both pipelined and end-to-end multi-turn interactions are inherently susceptible to error accumulation (Liao et al., 2018; 2021), it is crucial to validate whether intermediate outcomes consistently mediate between user inputs and agent outputs. In this context, the state consistency metric measures whether an agent correctly aligns user requests with its intermediate steps ($S_{t}$). Notably, agent’s intermediate steps can be represented in either a “structured format” (e.g., dialogue states) or “natural language” (e.g., Chain-of-Thought). By accommodating both representations, the state consistency can be formulated as follows:
$\text{{PIPA}}_{\mathcal{S}}=\frac{\displaystyle\sum_{t=1}^{T}\sum_{k=1}^{M_{t}}\text{IsConsistent}\bigl{(}S_{t}^{k},\;[\{U_{\tau},A_{\tau}\}_{\tau=1}^{t-1};U_{t}]\bigr{)}}{\displaystyle\sum_{t=1}^{T}M_{t}}$ (1)
Here, $S_{t}^{k}$ denotes the $k$-th state within the turn $t$. In addition, the IsConsistent function serves as a boolean indicator evaluated by LLMs that compares user requests (mostly natural language) with internal states (structured or natural language) [^2] Practically, as observed by Lee et al. (2025), evaluating a group of targeted samples collectively, using simple and atomic criteria such as boolean assessments, does not adversely affect performance compared to stepwise evaluation. Therefore, we group $S_{t}^{1:k}$ in the implementation..
@@ 3.1.2 Tool Efficiency ($\mathcal{A}$)
The management of external tools or functions, typically in the form of APIs, incurs operational costs and affects the task performance, underscoring the importance of evaluating their effective utilization. In particular, the occurrence of failed API calls prior to collecting complete information from users–commonly observed in multi-turn interactions–can result in unnecessary expenditures of resources (i.e., time and money) for service providers. To address this, the tool efficiency metric includes the total number of tool calls ($C_{T}$). For a given task, agent scenarios that require fewer API calls are considered indicative of more efficient API management.
Additionally, due to the stochastic nature of natural language prompting, generating well-structured API calls poses a significant challenge for certain agents. Even when an agent invokes an API at an appropriate time, iterative attempts at call generation will eventually lead to additional, avoidable costs. Consequently, this metric also calculates the number of failed tool generation attempts ($C_{F}$). Formally, we propose the following tool efficiency metric:
$\text{{PIPA}}_{\mathcal{A}}=\frac{C_{T}-C_{F}}{C_{T}+C_{F}}$ (2)
where the numerator rewards successful calls and the denominator penalizes excessive or failed calls. A higher value of $\text{{PIPA}}_{\mathcal{A}}$ indicates more efficient tool usage: maximizing successful calls while minimizing total and failed calls.
@@ 3.1.3 Observation Alignment ($\mathcal{O}$)
According to the original categorization of “qualitative measures” in Figure 2, criteria such as inappropriate response were used to indicate when the system’s output was factually, contextually, or semantically at odds with user’s needs (Walker et al., 1997). While such qualitative assessments provide valuable interpretative insights, they are difficult to incorporate directly into a formal decision‐theoretic framework. Consequently, we transition from a purely qualitative to a more quantitative representation so that the measure can be more easily integrated with our reward-based evaluation [^3] Crucially, this shift does not abandon the interpretive depth of the original qualitative measures; rather, it refines them into a metric that preserves underlying meanings and enables consistent, systematic evaluation across different dialogues and use cases..
On top of this discourse, the observation alignment metric evaluates whether observations appearing within the agent responses align with what the user requires. Specifically, it is calculated with a boolean criteria for each observation (i.e., retrieved database entity), preceded by capturing the number of observations ($|O|$) explicitly appear within the agent responses ($R_{agent}$). Considering the fact that both user utterance and agent response are represented in natural languages, it can be measured via a set of atomic LLM evaluations:
$\text{{PIPA}}_{\mathcal{O}}=\frac{1}{|T_{\text{obs}}|}\sum_{t\in T_{\text{obs}}}\left(\frac{1}{|O^{(t)}|}\sum_{o\in O^{(t)}}\text{IsAligned}(o,[\{U_{\tau},A_{\tau}\}_{\tau=1}^{t-1};U_{t}])\right)\,,$
$\text{where}\;T_{obs}=\{t\in T:|O^{(t)}|>0\}$ (3)
@@
3.1.4 Optional: Policy Alignment ($\mathcal{P}$)
Some interactive benchmarks assume certain policies (a behavioral rule appearing in agent responses, not only emerging in their internal states) that are globally reflected across the interactive sessions, such as Xiao et al. (2024). While some agents fail to provide tasks due to non-compliance with these policies, there are cases where agents do not follow the policy but succeed in achieving the intended goal (e.g., the policy “not recommending more than three items at a time” does not directly affect the task performance). Policy alignment is a session-level metric measuring whether a predefined set of policies ($\mathcal{P}$) are consistently followed throughout interactive sessions.
$\text{{PIPA}}_{\mathcal{P}}=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}{\text{IsAdherent}(p;\{U_{t},A_{t}\}_{t=1}^{T})}$ (4)
Similar to the state consistency metric (Section 3.1.1), IsAdherent serves as a boolean indicator deciding consistency between each policy and interaction context.

### v2

# AURA: A Diagnostic Framework for Tracking User Satisfaction of Interactive Planning Agents

@@ Abstract
The growing capabilities of large language models (LLMs) in instruction-following and context-understanding lead to the era of agents with numerous applications. Among these, task planning agents have become especially prominent in realistic scenarios involving complex internal pipelines, such as context understanding, tool management, and response generation. However, existing benchmarks predominantly evaluate agent performance based on task completion as a proxy for overall effectiveness. We hypothesize that merely improving task completion is misaligned with maximizing user satisfaction, as users interact with the entire agentic process and not only the end result. To address this gap, we propose AURA, an Agent-User inteRaction Assessment framework that conceptualizes the behavioral stages of interactive task planning agents. AURA offers a comprehensive assessment of agent through a set of atomic LLM evaluation criteria, allowing researchers and practitioners to diagnose specific strengths and weaknesses within the agent’s decision-making pipeline. Our analyses show that agents excel in different behavioral stages, with user satisfaction shaped by both outcomes and intermediate behaviors. We also highlight future directions, including systems that leverage multiple agents and the limitations of user simulators in task planning.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed in real-world applications, primarily due to their ability to understand complex goals and devise structured sequences of action: a process known as “planning” Wang_2024. To assess and refine planning skills, researchers have introduced diverse benchmarks in web-based NEURIPS2022_82ad13ec; xie2024openagents, mobile (deng-etal-2024-mobile), embodied (choi2024lotabench), and automated testbeds (Park2023GenerativeAgents).
One of the most impactful planning applications is task planning, where agents generate and execute domain-specific plans (e.g., itineraries) to help users achieve goals while adhering to contextual constraints (budzianowski-etal-2018-multiwoz; Rastogi_Zang_Sunkara_Gupta_Khaitan_2020; pmlr-v235-xie24j). While agents must accurately interpret user requests, leverage predefined tools, and engage in personalized dialogues, many benchmarks focus solely on final task completion, overlooking the agents’ intermediate behaviors and planning steps. Given that overall user satisfaction is shaped by multiple factors (e.g., efficiency and effectiveness) throughout prolonged interactions (ASHFAQ2020101473), a narrow focus on outcome-based metrics can misrepresent an agent’s true effectiveness, especially in complex task-oriented applications (cao2025revisitingllmevaluationmechanism; xia2025evaluationdrivendevelopmentllmagents; acikgoz-etal-2025-td).
In addition, existing benchmarks across different domains often introduce bespoke evaluation frameworks and metrics, making it possible for the same agent to be evaluated using entirely different criteria. This fragmentation introduces inconsistencies in how agent behavior is interpreted and makes it difficult to draw generalizable conclusions about agent capabilities. Without a domain-agnostic evaluation framework, it becomes challenging to meaningfully compare systems, track progress over time, and identify fundamental limitations in agent design.
To address these limitations, we propose AURA, an Agent-User inteRaction Assessment framework. To ensure generalizability across sequential agentic scenarios, AURA is designed based on the partially observed Markov Decision Process (POMDP). This formulation reflects the reality that key aspects of interaction, such as user intention and satisfaction, are often hidden or only partially observable. Moreover, it aligns with the observation that contemporary agent behaviors typically unfold in a sequential, decision-making context consistent with the POMDP paradigm (10610981; 10.5555/3692070.3692783; pmlr-v235-xie24j; yao2025taubench). Lastly, AURA defines domain-agnostic LLM evaluation criteria that can be easily instantiated through a set of atomic and easily measurable metrics. As illustrated in Figure 1, AURA provides the following key advantages:
Figure 1: AURA provides unified, atomic, and domain-agnostic criteria for assessing user satisfaction of interactive planning agents, extending beyond conventional evaluation protocols that focus solely on task completion.
@@
    Generalizable Evaluation: AURA establishes a domain-agnostic protocol for interactive task planning, enabling a consistent comparison across diverse benchmarks under a shared set of principles, which has been discussed for decades (walker-etal-1997-paradise) yet remains an open issue.
@@
    Multi-Axis Diagnosis: By aligning with the POMDP framework, AURA supports evaluation across multiple behavioral stages of the interaction pipeline, encompassing both intermediate decisions and final outcomes. This allows for a fine-grained diagnostic analysis of how different agents behave and impact performance.
@@
    Cross-Benchmark Comparisons: AURA enables systematic comparisons across heterogeneous tasks and environments, thereby uncovering planning strategies and revealing performance trade-offs that may not be observable within the scope of individual benchmarks.
Through extensive experiments, we demonstrate that different models exhibit distinct strengths and weaknesses within decision-making pipelines. Furthermore, human studies indicate that stage-specific evaluation using AURA correlates more strongly with improvements in user satisfaction than with the final task completion metric alone. Lastly, analyses on combining agents during deployment and the reliability of user simulators highlight promising research directions for future scholars and practitioners.
@@ 2 Background: Bespoke Evaluation of Planning Agents
Recent benchmarks for evaluating agentic tasks adopt rigid, domain-specific metrics to reflect the unique demands of each task setting, as summarized in Appendix A. This specialization is often necessary: interactive planning tasks vary widely in structure, goals, and interaction modalities, requiring tailored criteria to capture meaningful performance signals.
For instance, AgentBench (liu2024agentbench) incorporates domain-specific metrics such as Success Rate, Win Rate, F1 Score, and Exact Match to assess agents comprehensively. FlowBench (xiao-etal-2024-flowbench) further enhances the evaluation landscape by adding metrics focused explicitly on tool usage: Tool Invocation, measured by precision, recall, and F1 score for identifying the correct tool configurations; Success Rate, denoting the proportion of entirely successful sessions; and Task Progress, capturing the percentage of goals completed within a session. Meanwhile, $\tau$-Bench (yao2025taubench) utilizes more direct measures of success with Pass@k, a boolean indicator reflecting success across k attempts, and Pass^k, which requires successful outcomes for all k trials.
These metrics are often rigid by design, optimized for narrowly defined outcomes within their respective domains. However, this tight coupling limits their generalizability: it is difficult to apply these metrics across tasks or to capture broader notions of agent capability, especially in open-ended or compositional settings. As noted in prior work (ribeiro-etal-2020-beyond), overemphasis on final task success obscures nuanced failures or partial progress, particularly problematic in subjective or exploratory tasks like information aggregation (gangi-reddy-etal-2025-infogent; yoran-etal-2024-assistantbench). In contrast, AURA seeks to define a more general and domain-agnostic evaluation paradigm that is not tied to narrow success criteria.
3 Method
@@
Figure 2: Decision theory-based taxonomy of evaluation metrics in AURA. As discussed in walker-etal-1997-paradise, it should be cautious to generalize the original taxonomy to different agents and tasks. Following this, we provide five distinct interpretations of each element, which will be described in Section 3.2.
3.1 Metric Design Criteria
Interactive planning tasks are generally modeled using the POMDP paradigm, where agents devise plans based on partial observational data (10610981; 10.5555/3692070.3692783; pmlr-v235-xie24j; yao2025taubench). Building on the framework of WILLIAMS2007393, we present an abstract and discrete pipeline consisting of $\mathcal{S}$ (a set of agent states), $\mathcal{A}$ (a set of actions), $\mathcal{O}$ (a set of observations) and $\mathcal{R}$ (a set of rewards), along with an our additional policy definition $\mathcal{P}$ (a set of global policies). While task completion is a common performance goal, we hypothesize that it is insufficient: user satisfaction depends on how an agent arrives at outcomes (ASHFAQ2020101473; 10.1145/3477495.3531798).
To that end, AURA decomposes evaluation into atomic, phase-specific criteria reflecting the multi-stage behavior of LLM agents. These include state consistency, tool efficiency, observation alignment, policy adherence, and task completion, each tied to a specific step in the decision pipeline. Drawing from prior work on decomposed evaluation (liu-etal-2023-revisiting; saadfalcon2024lmunitfinegrainedevaluationnatural; lee2024checkevalrobustevaluationframework), we design intuitive LLM-based evaluations for each metric for more consistent and precise diagnostics.
Furthermore, as illustrated in Figure 2, we align AURA with the decision theoretic principles (walker-etal-1997-paradise), where disparate metrics collectively estimate user satisfaction. POMDPs provide a natural foundation here, as agents select actions under uncertainty to maximize expected utility, capturing both goal achievement and procedural quality. This integration enables AURA to measure user satisfaction holistically while remaining grounded in theoretical rigor.
Throughout this work, we formulate the basic components of multi-turn, multi-step agents as follows: $T$ denotes the total number of turns (i.e., alternating utterances between the user and the agent) in an interactive session, $M_{t}$ represents the number of steps in the agent’s internal reasoning process within a single utterance at turn $t$, and $\mathcal{C}_{t}=\{u_{\tau},a_{\tau}\}_{\tau=1}^{t}$ signifies the context of user inputs and agent responses observed up to turn $t$.
3.2 Components of AURA
@@ 3.2.1 State Consistency ($\mathcal{S}$)
The intermediate decision-making step of the agent is pivotal in summarizing user requests and determining optimal subsequent actions (shin-etal-2022-dialogue; wei2022chain). Given that both pipelined and end-to-end multi-turn interactions are inherently susceptible to error accumulation (10.1145/3240508.3240605; liao-etal-2021-dialogue), it is crucial to validate whether intermediate outcomes consistently mediate between user inputs and agent outputs. In this context, the state consistency metric measures whether an agent correctly aligns user requests with its $k$-th intermediate steps ($z_{t}^{k}$). Notably, these intermediate steps can be represented in either a “structured format” (e.g., dialogue states) or “natural language” (e.g., Chain-of-Thought). By accommodating both representations, the state consistency can be formulated as follows:
[…]
Here, the IsConsistent function serves as a boolean indicator evaluated by LLMs that compares user requests (mostly natural language) with internal states (structured or natural language) [^1] Practically, as observed by lee2024checkevalrobustevaluationframework, evaluating a group of targeted samples collectively, using simple and atomic criteria such as boolean assessments, does not adversely affect performance compared to stepwise evaluation. Therefore, we group $z_{t}^{1:k}$ in the implementation..
@@ 3.2.2 Tool Efficiency ($\mathcal{A}$)
The management of external tools or functions, typically in the form of APIs, incurs operational costs and affects the task performance, underscoring the importance of evaluating their effective utilization. In particular, the occurrence of failed API calls prior to collecting complete information from users can result in unnecessary expenditures of resources (i.e., time and money) and negative user experience. To address this, the tool efficiency metric first considers the total number of tool calls ($N_{T}$). For a given task, agent scenarios that require fewer API calls are considered indicative of more efficient API management.
Additionally, due to the stochastic nature of natural language prompting, generating well-structured API calls poses a significant challenge for certain agents. Even when an agent invokes an API at an appropriate time, iterative attempts due to incomplete generation will eventually lead to additional, avoidable costs. Consequently, this metric also calculates the number of failed tool generation attempts ($N_{F}$). Formally, we propose the following tool efficiency measure:
$\text{{AURA}}_{\mathcal{A}}=\frac{N_{T}-N_{F}}{N_{T}+N_{F}}$
where the numerator rewards successful calls and the denominator penalizes excessive or failed calls. A higher value of $\text{{AURA}}_{\mathcal{A}}$ indicates more efficient tool usage: maximizing successful calls while minimizing total and failed calls.
@@ 3.2.3 Observation Alignment ($\mathcal{O}$)
The observation alignment metric evaluates whether observations appearing within the context align with what the user requires. Specifically, it is calculated with a boolean criteria for each observation (i.e., retrieved database entity), preceded by capturing the number of observations ($|O|$) explicitly appear within the agent responses. Considering the fact that both user utterance and agent response are represented in natural languages, it can be measured via a set of atomic LLM evaluations:
$\text{{AURA}}_{\mathcal{O}}=\frac{1}{|T_{\text{obs}}|}\sum_{t\in T_{\text{obs}}}\left(\frac{1}{|O^{(t)}|}\sum_{o\in O^{(t)}}\text{IsAligned}(o,\mathcal{C}_{t}\backslash\{a_{t}\})\right)\,,\text{where}\;T_{obs}=\{t\in T:|O^{(t)}|>0\}$
@@
3.2.4 Policy Alignment ($\mathcal{P}$)
Interactive agent benchmarks assume certain policies (i.e., a behavioral rule appearing in agent responses, not only emerging in their internal states) that are globally reflected across the interactive sessions (xiao-etal-2024-flowbench), mostly in a form of system prompt.
The Policy alignment is a session-level metric measuring whether a predefined set of policies ($\mathcal{P}$) are consistently followed throughout interactive sessions.
$\text{{AURA}}_{\mathcal{P}}=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}{\text{IsAdherent}(p;\mathcal{C}_{T})}$
Similar to the state consistency (Section 3.2.1) and observation alignment (Section 3.2.3) metrics, IsAdherent serves as an LLM-evaluated boolean indicator deciding consistency between each policy and interaction context.
# of Scenarios # of Tools # of Database
TravelPlanner pmlr-v235-xie24j 180 7 3,865,195 total, 3,827,361 max for a tool
$\tau$-Bench-Airline yao2025taubench 50 13 500 users, 300 flights, 2,000 reservations
$\tau$-Bench-Retail yao2025taubench 115 15 500 users, 50 products, 1,000 orders